OpenStudy (anonymous):

Forgot how to do this: Find the circumference of a circle if C=7.76 and the n thing is 3.14

OpenStudy (anonymous):

C=2nr Don't know how to make that little thing whatever you call it.

OpenStudy (anonymous):

It's called pi. 

OpenStudy (anonymous):

Now, divide both side by 2n.

OpenStudy (anonymous):

194 over 157?

OpenStudy (anonymous):

Wait, by C, you mean radius, right?

OpenStudy (anonymous):

It says to solve for the missing variable and the blank where the answer goes is r=

OpenStudy (anonymous):

In the question, you asked us to find the circumference of a circle. Do you actually mean radius? (r is radius)

OpenStudy (anonymous):

I don't know the question that I got it from just says this exactly:

OpenStudy (anonymous):

That is the picture of the actual question.

OpenStudy (anonymous):

Yeah, it asking for radius. We know that C = 7.76 and 2n = 6.28 \[C = 2{\pi}r \rightarrow 7.76 = 6.28r \rightarrow \frac{ 7.76 }{ 6.28 } = \frac{ 6.28 r }{ 6.28 }\]\[\frac{ 7.76 }{ 6.28 } = r\] Now, calculate it.

OpenStudy (anonymous):

I keep getting 194/157.